# A $2$-distance $(2\Delta+7)$-coloring of planar graphs

### v2

@@ 2.1. The case $\Delta=6$
Let $v$ be a $5(4)$-vertex and denote by $f_{1},f_{2},\ldots,f_{5}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$.
$(a)$ Let $m_{4}(v)=0$. Assume for a contradiction that $v$ has at most one $6(5^{-})$-neighbour. In this case, $v$ has four neighbours consisting of 6(6)666(6)6 ( 6 )- or $5^{-}$-vertices, and so we have $d_{2}(v)\leq 18$. If we set $G^{\prime}=G-v+\{v_{1}v_{5}\}$, then $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction.
$(b)$ Let $m_{4}(v)=1$. By contradiction, assume that $v$ has at most two $6(5^{-})$-neighbours. In this case, $v$ has three neighbours consisting of 6(6)666(6)6 ( 6 )- or $5^{-}$-vertices, and so $d_{2}(v)\leq 18$. Similarly as above, we get a contradiction. ∎
@@
Let $v$ be a $5(4)$-vertex and denote by $f_{1},f_{2},\ldots,f_{5}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$.
Assume that $v$ has two $5(4)$-neighbours $v_{i},v_{j}$ such that $|i-j|\geq 2$. Then $v_{i}v_{i-1}$ or $v_{i}v_{i+1}$ (in a cyclic ordering) is contained in two $3$-faces. Moreover, $v_{j}v_{j-1}$ or $v_{j}v_{j+1}$ (in a cyclic ordering) is contained in two $3$-faces. We then deduce that $d_{2}(v)\leq 18$. If we set $G^{\prime}=G-v+\{v_{1}v_{5}\}$, then $G^{\prime}$ would be proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction. ∎
@@
Let $v$ be a 5(5)555(5)5 ( 5 )-vertex. Then $v$ has neither $4$-neighbour nor bad $5$-neighbour nor 6(6)666(6)6 ( 6 )-neighbour. In particular, $n_{5}(v)\leq 1$.
@@
Let $v$ be a 5(5)555(5)5 ( 5 )-vertex and denote by $f_{1},f_{2},\ldots,f_{5}$ the $3$-faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$ in a cyclic fashion.
Recall that $G$ has no $4(3^{+})$-vertex by Lemma 2.3, also $v$ cannot have any $4(1^{-})$-neighbour as $m_{3}(v)=5$. Moreover, $v$ cannot have any $4(2)$-neighbour by Lemma 2.7. Consequently, we deduce that $v$ has no $4$-neighbour. On the other hand, if $v$ has a bad $5$-neighbour or a 6(6)666(6)6 ( 6 )-neighbour or two $5$-neighbours, then we have $d_{2}(v)\leq 18$. When we set $G^{\prime}=G-v$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction. ∎
The following is an easy consequence of Lemma 2.12.
@@
If $v$ is a 5(5)555(5)5 ( 5 )-vertex, then $v$ has four $6(5^{-})$-neighbours.
An edge $uv$ is said to be special if $v$ is a 5(5)555(5)5 ( 5 )-vertex, and $uv$ is contained in two $3$-faces $f_{1},f_{2}$ such that each of those faces is adjacent to a $4^{+}$-face (see Figure 1). Recall that $u$ is a $4^{+}$-vertex by Lemma 2.2.
@@
Figure 1. A special edge $uv$ with a 5(5)555(5)5 ( 5 )-vertex $v$ and a $4^{+}$-vertex $u$.
@@
    If $uw$ is contained in two $3$-faces, then all neighbours of $v$ are $6(5^{-})$-vertices. In particular, for every vertex $z\in N(v)\setminus\{u,w\}$, the edge $vz$ is a special edge.
@@
    If no edge in $G[N(v)]$ is contained in two $3$-faces, then $v$ has four $6(4^{-})$-neighbours, and each $vv_{i}$ is a special edge for $v_{i}\in N(v)$.
@@
Let $v$ be a 5(5)555(5)5 ( 5 )-vertex, and denote by $f_{1},f_{2},\ldots,f_{5}$ the $3$-faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$ in a cyclic fashion.
$(a)$ Suppose first that there exists $v_{i}\in N(v)$ such that $v_{i}v_{i+1}$ is contained in two $3$-faces. If $v$ has a $5^{-}$- or 6(6)666(6)6 ( 6 )-neighbours, then $d_{2}(v)\leq 18$. When we set $G^{\prime}=G-v$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction. Thus, all neighbours of $v$ are $6(5^{-})$-vertices. In particular, if there also exists $v_{j}\in N(v)\setminus\{v_{i},v_{i+1}\}$ such that $v_{j}v_{j-1}$ or $v_{j}v_{j+1}$ is contained in two $3$-faces, then we have again $d_{2}(v)\leq 18$. Similarly as above, we get a contradiction. Therefore, for every vertex $v_{j}\in N(v)\setminus\{v_{i},v_{i+1}\}$, the edge $vv_{j}$ is a special edge.
$(b)$ Suppose that no edge in $G[N(v)]$ is contained in two $3$-faces. This implies that each $vv_{i}$ is a special edge for $v_{i}\in N(v)$. So, the claim follows from Corollary 2.13. ∎
@@
Let $v$ be a $6(5)$-vertex. Then $n_{4}(v)\leq 3$. In particular,
@@
    if $n_{4}(v)=2$, then $v$ has at most two bad $5$-neighbours,
@@
    if $n_{4}(v)=2$ and $v$ has two bad $5$-neighbours, then $v$ has two $6(4^{-})$-neighbours,
@@
    if $n_{4}(v)=1$, then $v$ has at most three bad $5$-neighbours,
@@
    if $n_{4}(v)=1$ and $v$ has three bad $5$-neighbours, then $v$ has a $6(4^{-})$-neighbour,
@@
    if $n_{4}(v)=0$, then $v$ has at most four bad $5$-neighbours,
@@
    if $n_{4}(v)=0$, $m_{4}(v)=1$ and $v$ has four bad $5$-neighbours, then $v$ has two $6(4^{-})$-neighbours.
@@
Let $v$ be a $6(5)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{6}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$. Recall first that $v$ has no $3$-neighbour by Lemma 2.2.
We now claim that $n_{4}(v)\leq 3$. Otherwise, $v$ has four $4$-neighbours, and so $d_{2}(v)\leq 18$. By assumption, one of $v_{2},v_{3},v_{4}$ must be $4$-vertex, say $v_{2}$. If we set $G^{\prime}=G-v+\{v_{2}v_{4},v_{2}v_{5},v_{2}v_{6}\}$, then $G^{\prime}$ would be proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $19$ colors. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction. Hence, $n_{4}(v)\leq 3$.
$(a)$ Let $n_{4}(v)=3$. By contradiction, assume that $v$ has a bad $5$-neighbour, say $v_{i}$. This forces that either $v_{i}v_{i-1}$ or $v_{i}v_{i+1}$ is contained in two $3$-faces. In each case, we have $d_{2}(v)\leq 18$. By assumption, one of $v_{2},v_{3},v_{4},v_{5}$ is a $4$-vertex, say $v_{2}$. Let $G^{\prime}=G-v+\{v_{2}v_{4},v_{2}v_{5},v_{2}v_{6}\}$. Clearly, $G^{\prime}$ is proper with respect to $G$. Similarly as above, we get a contradiction.
@@
$(b)$ Let $n_{4}(v)=2$. Assume to the contrary that $v$ has three bad $5$-neighbours. This forces that there exist two distinct vertices $v_{i},v_{j}\in N(v)$ such that $v_{i}v_{i+1}$ and $v_{j}v_{j+1}$ are contained in two $3$-faces. It is clear that $d_{2}(v)\leq 17$. By assumption, $v$ has two $4$-neighbours. If one of $v_{2},v_{3},v_{4},v_{5}$ is a $4$-vertex, say $v_{2}$, then we set $G^{\prime}=G-v+\{v_{2}v_{4},v_{2}v_{5},v_{2}v_{6}\}$, and so $G^{\prime}$ would be proper with respect to $G$. Otherwise, suppose that both $v_{1}$ and $v_{6}$ are $4$-vertices. Consider a $5$-neighbour of $v$, say $v_{2}$, if we set $G^{\prime}=G-v+\{v_{1}v_{6},v_{2}v_{4},v_{2}v_{6}\}$, then $G^{\prime}$ would be proper with respect to $G$. Similarly as above, we get a contradiction.
$(c)$ Let $n_{4}(v)=2$, and suppose that $v$ has two bad $5$-neighbours, say $v_{i},v_{j}$. This forces that either $v_{i}v_{i-1}$ or $v_{i}v_{i+1}$ is contained in two $3$-faces. Moreover, if $v_{i}$ and $v_{j}$ are non-adjacent, then there exists another edge forming $v_{j}v_{j-1}$ or $v_{j}v_{j+1}$ such that it is contained in two $3$-faces. We remark that if $v$ has at most one $6(4^{-})$-neighbour, then the last neighbour of $v$ other than $v_{i},v_{j}$, $4$-vertex and $6(4^{-})$-vertex would be either $5$-vertex or a $6(5)$-vertex or a 6(6)666(6)6 ( 6 )-vertex. In each case, we have $d_{2}(v)\leq 18$. By applying a similar process to $(b)$, we get a contradiction.
@@
$(d)$ Let $n_{4}(v)=1$. Assume that $v$ has four bad $5$-neighbours. This forces that there exist two distinct vertices $v_{i},v_{j}\in N(v)$ with $|i-j|>1$ such that $v_{i}v_{i+1}$ and $v_{j}v_{j+1}$ are contained in two $3$-faces. Clearly we have $d_{2}(v)\leq 18$. By assumption, $v$ has a $4$-vertex. If one of $v_{2},v_{3},v_{4},v_{5}$ is a $4$-vertex, say $v_{2}$, then we set $G^{\prime}=G-v+\{v_{2}v_{4},v_{2}v_{5},v_{2}v_{6}\}$, and so $G^{\prime}$ would be proper with respect to $G$. Otherwise, suppose that $v_{1}$ (or $v_{6}$) is a $4$-vertex. Since $v$ has four $5$-neighbours, one of $v_{3},v_{5}$ must be $5$-vertex, say $v_{3}$. Let $G^{\prime}=G-v+\{v_{1}v_{6},v_{3}v_{1},v_{3}v_{5}\}$. Clearly, $G^{\prime}$ is proper with respect to $G$. Similarly as above, we get a contradiction.
$(e)$ Let $n_{4}(v)=1$, and suppose that $v$ has three bad $5$-neighbours, say $v_{i},v_{j},v_{k}$. This forces that there exist two distinct vertices $v_{p},v_{q}\in N(v)$ such that $v_{p}v_{p+1}$ and $v_{q}v_{q+1}$ are contained in two $3$-faces. If $v$ has no $6(4^{-})$-neighbour, then the last two neighbours of $v$ other than $v_{i},v_{j},v_{k}$ and $4$-vertex consist of $5$-, $6(5)$- or 6(6)666(6)6 ( 6 )-vertices. In each case, we have $d_{2}(v)\leq 18$. Similar to $(d)$, we get a contradiction.
@@
$(f)$ Let $n_{4}(v)=0$. Assume to the contrary that $v$ has five bad $5$-neighbours. This forces that there exist three distinct vertices $v_{i},v_{j},v_{t}\in N(v)$ such that each of $v_{i}v_{i+1}$, $v_{j}v_{j+1}$, $v_{t}v_{t+1}$ is contained in two $3$-faces. Clearly, we have $d_{2}(v)\leq 18$. Since $v$ has five bad $5$-neighbours, one of $v_{1},v_{6}$ is a $5$-vertex, say $v_{1}$. Moreover, one of $v_{3},v_{5}$ is a $5$-vertex, say $v_{3}$. Let $G^{\prime}=G-v+\{v_{1}v_{6},v_{3}v_{1},v_{3}v_{5}\}$. Clearly, $G^{\prime}$ is proper with respect to $G$. Similarly as above, we get a contradiction.
$(g)$ Suppose that $n_{4}(v)=0$, $m_{4}(v)=1$, and $v$ has four bad $5$-neighbours. Since $v$ has four bad $5$-neighbours, there exist two distinct vertices $v_{i},v_{j}\in N(v)$ with $|i-j|>1$ such that $v_{i}v_{i+1}$ and $v_{j}v_{j+1}$ are contained in two $3$-faces. By contradiction, assume that $v$ has at most one $6(4^{-})$-neighbour. Then the sixth neighbour of $v$ other than bad $5$-vertex and $6(4^{-})$-vertex is either a $5(t)$-vertex with $t\leq 3$ or a $6(5^{+})$-vertex. Clearly, we have $d_{2}(v)\leq 18$. Moreover, this particularly implies that either $v$ has totally five $5$-neighbours or there exist three distinct vertices $v_{i},v_{j},v_{t}\in N(v)$ such that each of $v_{i}v_{i+1}$, $v_{j}v_{j+1}$, $v_{t}v_{t+1}$ is contained in two $3$-faces. If $v$ has five $5$-neighbours, then one of $v_{1},v_{6}$ must be a $5$-vertex, say $v_{1}$. Also, one of $v_{3},v_{5}$ must be a $5$-vertex, say $v_{3}$. Let $G^{\prime}=G-v+\{v_{1}v_{6},v_{3}v_{1},v_{3}v_{5}\}$. Clearly, $G^{\prime}$ is proper with respect to $G$. Similarly as above, we get a contradiction.
@@
We now suppose that $v$ has exactly four $5$-neighbours. If $v_{1}$ (resp. $v_{6}$) is a $5$-vertex, and one of $v_{3},v_{5}$ (resp. $v_{2},v_{4}$) is a $5$-vertex, then we apply the same argument as above. Therefore, we may assume that $v_{1},v_{6}$ are $6$-vertices and $v_{2},v_{3},v_{4},v_{5}$ are bad $5$-vertices. In fact, $v_{2},v_{3},v_{4},v_{5}$ must be $5(4)$-vertices, since a 5(5)555(5)5 ( 5 )-vertex cannot have any bad $5$-neighbour by Lemma 2.12. However, this yields a contradiction by Proposition 2.11. ∎
@@
We will now apply discharging to show that $G$ does not exist. First, we assign to each vertex $v$ a charge $\mu(v)=d(v)-4$ and to each face $f$ a charge $\mu(f)=\ell(f)-4$. By Euler’s formula, we have
@@
      $f$ gives $\frac{1}{3}$ to each of its incident $3$-vertices and bad $4$-vertices,
@@
      $f$ gives $\frac{1}{5}$ to each of its incident $5(4)$-vertices,
@@
      $f$ gives $\frac{1}{5}$ to each of its incident $6(5)$-vertex $u$ if $n_{4}(u)=0$.
@@
    Every $6(k)$-vertex for $0\leq k\leq 3$ gives $\frac{1}{6}$ to each of its neighbours.
@@
      $v$ gives $\frac{1}{12}$ to each of its $4$- and $6(5)$-neighbours,
@@
      $v$ gives $\frac{1}{9}$ to each of its $3$- and $5(4)$-neighbours,
@@
      $v$ gives $\frac{1}{6}$ to each of its 5(5)555(5)5 ( 5 )-neighbour $u$ if $uv$ is a special edge,
@@
      $v$ gives $\frac{1}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbour $u$ if $uv$ is not a special edge.
@@
      $v$ gives $\frac{1}{12}$ to each of its bad $4$-neighbours,
@@
      $v$ gives $\frac{1}{9}$ to each of its bad $5$-neighbours.

### v3

@@ 2.1. The case $\Delta=6$
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$.
$(a)$ Let $m_{4}(v)=0$ and assume, for a contradiction, that $v$ has at most one $6(5^{-})$-neighbour. Then $v$ has four neighbours that are all $6(6)$- or $5^{-}$-vertices, and hence $d_{2}(v)\leq 18$ by Lemma 2.2. If we define $G^{\prime}=G-v+\{v_{1}v_{5}\}$, then $G^{\prime}$ is proper with respect to $G$. By Remark 2.1, this yields a contradiction.
$(b)$ Let $m_{4}(v)=1$ and assume, again for a contradiction, that $v$ has at most two $6(5^{-})$-neighbours. Then $v$ has three neighbours that are $6(6)$- or $5^{-}$-vertices, implying $d_{2}(v)\leq 18$ by Lemma 2.2. As in the previous case, this leads to a contradiction. ∎
@@
Let $v$ be a $5(4)$-vertex, and let $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$. Assume that $v$ has two $5(4)$-neighbours $v_{i}$ and $v_{j}$ such that $|i-j|\geq 2$. Then either $v_{i}v_{i-1}$ or $v_{i}v_{i+1}$ (in cyclic order) is contained in two $3$-faces. Similarly, either $v_{j}v_{j-1}$ or $v_{j}v_{j+1}$ is contained in two $3$-faces. Thus $d_{2}(v)\leq 18$ by Lemma 2.2. Setting $G^{\prime}=G-v+\{v_{1}v_{5}\}$ yields a graph $G^{\prime}$ that is proper with respect to $G$, giving a contradiction by Remark 2.1. ∎
@@
Let $v$ be a $5(5)$-vertex. Then $v$ has neither a $4$-neighbour nor a bad $5$-neighbour nor a $6(6)$-neighbour. In particular, $n_{5}(v)\leq 1$.
@@
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$ in cyclic order. By Lemma 2.5, $v$ has no $4$-neighbour. Now suppose that $v$ has either a bad $5$-neighbour, or a $6(6)$-neighbour, or two $5$-neighbours. In each of these cases, Lemma 2.2 gives $d_{2}(v)\leq 18$. If we set $G^{\prime}=G-v$, then $G^{\prime}$ remains proper with respect to $G$. By Remark 2.1, this contradicts the minimality of $G$. ∎
The following is an immediate consequence of Lemma 2.14.
@@
If $v$ is a $5(5)$-vertex, then $v$ has four $6(5^{-})$-neighbours.
An edge $uv$ is said to be special if $v$ is a $5(5)$-vertex and $uv$ is contained in two $3$-faces $f_{1},f_{2}$, each of which is adjacent to a $4^{+}$-face (see Figure 1). Note that $u$ is a $5^{+}$-vertex by Lemmas 2.4 and 2.5,
@@
Figure 1. A special edge uvuv with a $5(5)$-vertex $v$ and a $5^{+}$-vertex $u$.
@@
    If $v_{i}v_{i+1}$ is contained in two $3$-faces, then all neighbours of $v$ are $6(5^{-})$-vertices. In particular, for every vertex $v_{j}\in N(v)\setminus\{v_{i},v_{i+1}\}$, the edge $vv_{j}$ is a special edge.
@@
    If no edge in $E(v)$ is contained in two $3$-faces, then each edge $vv_{i}$ is a special edge for every $v_{i}\in N(v)$, and $v$ has four $6(4^{-})$-neighbours.
@@
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$ in cyclic order.
$(a)$ Suppose that the edge $v_{i}v_{i+1}$ is contained in two $3$-faces. If $v$ has either a $5^{-}$-neighbour or a $6(6)$-neighbour, then Lemma 2.2 implies $d_{2}(v)\leq 18$. Removing $v$ (i.e., taking $G^{\prime}=G-v$) yields a graph proper with respect to $G$, contradicting Remark 2.1. Hence every neighbour of $v$ must be a $6(5^{-})$-vertex.
Now suppose that there exists another neighbour $v_{j}\in N(v)\setminus\{v_{i},v_{i+1}\}$ such that either $v_{j}v_{j-1}$ or $v_{j}v_{j+1}$ is contained in two $3$-faces. Then Lemma 2.2 again gives $d_{2}(v)\leq 18$, producing the same contradiction. Therefore, for every $v_{j}\in N(v)\setminus\{v_{i},v_{i+1}\}$, the edge $vv_{j}$ is a special edge.
$(b)$ Suppose no edge in $E(v)$ is contained in two $3$-faces. Then, by definition of a special edge, each $vv_{i}$ is a special edge for all $v_{i}\in N(v)$. Since a $5(5)$-vertex has no bad $5$-neighbour and no $6(6)$-neighbour by Lemma 2.14, it follows from Corollary 2.15 that $v$ has four $6(5^{-})$-neighbours. ∎
@@
Let $v$ be a $6(5)$-vertex. Then $n_{4}(v)\leq 3$. In particular:
@@
    If $n_{4}(v)=2$, then $v$ has at most two bad $5$-neighbours.
@@
    If $n_{4}(v)=2$ and $v$ has two bad $5$-neighbours, then $v$ has two $6(4^{-})$-neighbours.
@@
    If $n_{4}(v)=1$, then $v$ has at most three bad $5$-neighbours.
@@
    If $n_{4}(v)=1$ and $v$ has three bad $5$-neighbours, then $v$ has a $6(4^{-})$-neighbour.
@@
    If $n_{4}(v)=0$, then $v$ has at most four bad $5$-neighbours.
@@
    If $n_{4}(v)=0$, $m_{4}(v)=1$, and $v$ has four bad $5$-neighbours, then $v$ has two $6(4^{-})$-neighbours.
@@
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$. Recall that $v$ has no $3$-neighbour by Lemma 2.4.
We claim that $n_{4}(v)\leq 3$. Indeed, if $v$ had four $4$-neighbours, then Lemma 2.2 would imply $d_{2}(v)\leq 18$. Among $v_{2},v_{3},v_{4},v_{5}$, at least one must be a $4$-vertex; assume $v_{2}$ is such. Then setting $G^{\prime}=G-v+\{v_{2}v_{4},v_{2}v_{5},v_{2}v_{6}\}$ yields a graph that is proper with respect to $G$, contradicting Remark 2.1.
$(a)$ Let $n_{4}(v)=3$. If $v$ had a bad $5$-neighbour $v_{i}$, then by Lemma 2.2 we would have $d_{2}(v)\leq 18$. Similarly as above, this leads to a contradiction.
@@
$(b)$ Let $n_{4}(v)=2$. If $v$ had three bad $5$-neighbours, then by Lemma 2.2 we would have $d_{2}(v)\leq 17$. If one of $v_{2},v_{3},v_{4},v_{5}$ is a $4$-vertex (say $v_{2}$), take $G^{\prime}=G-v+\{v_{2}v_{4},v_{2}v_{5},v_{2}v_{6}\}$. Otherwise, both $v_{1}$ and $v_{6}$ are $4$-vertices; with a $5$-neighbour $v_{2}$, we may take $G^{\prime}=G-v+\{v_{1}v_{6},v_{2}v_{4},v_{2}v_{6}\}$. In both cases, $G^{\prime}$ is proper with respect to $G$, a contradiction.
$(c)$ Let $n_{4}(v)=2$, and suppose that $v$ has two bad $5$-neighbours. Each bad $5$-neighbour $v_{i}$ forces either $v_{i}v_{i-1}$ or $v_{i}v_{i+1}$ to lie in two $3$-faces. If $v$ had at most one $6(4^{-})$-neighbour, then the last neighbour of $v$ would be either a $5$- or $6(5)$- or $6(6)$-vertex, giving $d_{2}(v)\leq 18$. Applying the same reductions as in part (b) yields a contradiction. Thus $v$ has at least two $6(4^{-})$-neighbours.
@@
$(d)$ Let $n_{4}(v)=1$. If $v$ had four bad $5$-neighbours, then by Lemma 2.2 we would have $d_{2}(v)\leq 18$. If one of $v_{2},v_{3},v_{4},v_{5}$ is a $4$-vertex (say $v_{2}$), use $G^{\prime}=G-v+\{v_{2}v_{4},v_{2}v_{5},v_{2}v_{6}\}$. Otherwise the $4$-vertex is $v_{1}$ or $v_{6}$; assume $v_{1}$. Since $v$ has four $5$-neighbours, one of $v_{3},v_{5}$ is $5$; assume $v_{3}$. Then $G^{\prime}=G-v+\{v_{1}v_{6},v_{3}v_{1},v_{3}v_{5}\}$ is proper with respect to $G$, a contradiction.
$(e)$ Let $n_{4}(v)=1$ and suppose that $v$ has three bad $5$-neighbours. Observe that two distinct edges among $v_{i}v_{i+1}$ and $v_{j}v_{j+1}$ are contained in two $3$-faces. If $v$ had no $6(4^{-})$-neighbour, then its remaining neighbours would be $5$-, $6(5)$-, or $6(6)$-vertices, giving $d_{2}(v)\leq 18$ by Lemma 2.2. If one of $v_{2},v_{3},v_{4},v_{5}$ is a $4$-vertex (say $v_{2}$), then define $G^{\prime}=G-v+\{v_{2}v_{4},v_{2}v_{5},v_{2}v_{6}\}$. Otherwise, the $4$-vertex must be $v_{1}$ or $v_{6}$; assume $v_{1}$. If $v_{2},v_{4},v_{6}$ are $5$-vertices, then we set $G^{\prime}=G-v+\{v_{1}v_{6},v_{2}v_{4},v_{2}v_{6}\}$. Else, one of $v_{3},v_{5}$ is a $5$-vertex; assume $v_{3}$. Then $G^{\prime}=G-v+\{v_{1}v_{6},v_{3}v_{1},v_{3}v_{5}\}$. In each case, $G^{\prime}$ is proper with respect to $G$, contradicting Remark 2.1. Therefore, $v$ must have a $6(4^{-})$-neighbour.
@@
$(f)$ Let $n_{4}(v)=0$. If $v$ had five bad $5$-neighbours, then there would exist three distinct edges of $E(v)$ such that each of them is contained in two $3$-faces, giving $d_{2}(v)\leq 18$ by Lemma 2.2. Among $v_{1},v_{6}$, one is $5$-vertex; assume $v_{1}$. Among $v_{3},v_{5}$, one is $5$-vertex; assume $v_{3}$. Then $G^{\prime}=G-v+\{v_{1}v_{6},v_{3}v_{1},v_{3}v_{5}\}$ is proper with respect to $G$, a contradiction. Thus $v$ has at most four bad $5$-neighbours.
$(g)$ Let $n_{4}(v)=0$, $m_{4}(v)=1$, and suppose that $v$ has four bad $5$-neighbours. Observe that two non-adjacent edges among $v_{i}v_{i+1}$ is contained in two $3$-faces. If $v$ had at most one $6(4^{-})$-neighbour, then its remaining neighbour would be a $5(3^{-})$-vertex or a $6(5^{+})$-vertex, giving $d_{2}(v)\leq 18$ by Lemma 2.2. This implies either $v$ has five $5$-neighbours or three edges around $v$ lie in two $3$-faces.
@@
If $v$ has five $5$-neighbours, the argument from part (f) applies, giving a contradiction. Thus $v$ has exactly four $5$-neighbours. If $v_{1}$ (resp. $v_{6}$) is a $5$-vertex and one of $v_{3},v_{5}$ (resp. $v_{2},v_{4}$) is a $5$-vertex, we again use the same reduction as in (f). Therefore we may assume $v_{1},v_{6}$ are $6$-vertices and $v_{2},v_{3},v_{4},v_{5}$ are bad $5$-vertices. Since a $5(5)$-vertex cannot have any bad $5$-neighbour by Lemma 2.14, each of $v_{2},v_{3},v_{4},v_{5}$ must be a $5(4)$-vertex. However, by Proposition 2.13, a $5(4)$-vertex cannot have two non-adjacent $5(4)$-neighbours. This contradicts planarity of $G$, and therefore $v$ must have at least two $6(4^{-})$-neighbours. ∎
@@
We now apply discharging method to show that $G$ cannot exist. First, we assign to each vertex $v$ a charge $\mu(v)=d(v)-4$ and to each face $f$ a charge $\mu(f)=\ell(f)-4$. By Euler’s formula, we have
@@
      $\frac{1}{3}$ to each of its incident $3$-vertices and bad $4$-vertices,
@@
      $\frac{1}{5}$ to each of its incident $5(4)$-vertices,
@@
      $\frac{1}{5}$ to each of its incident $6(5)$-vertex having no $4$-neighbour.
@@
    Every $6(3^{-})$-vertex gives $\frac{1}{6}$ to each of its neighbours.
@@
      $\frac{1}{12}$ to each of its $4$- and $6(5)$-neighbours,
@@
      $\frac{1}{9}$ to each of its $3$- and $5(4)$-neighbours,
@@
      $\frac{1}{6}$ to each of its $5(5)$-neighbour $u$ if $uv$ is a special edge,
@@
      $\frac{1}{9}$ to each of its $5(5)$-neighbour $u$ if $uv$ is not a special edge.
@@
      $\frac{1}{12}$ to each of its bad $4$-neighbours,
@@
      $\frac{1}{9}$ to each of its bad $5$-neighbours.